### v1

# New Skyrme parametrizations to describe finite nuclei and neutron star matter with realistic effective masses. II. Adjusting the spin-dependent terms

Abstract
Many common Skyrme functionals present ferromagnetic instabilities or unrealistic density dependence of the spin-1 Landau parameters. To solve these problems, we consider the Skyrme interaction as a density-functional rather than a density-dependent two-body force. This allows us to adjust the spin-dependent terms of the new extended Skyrme functionals of our previous paper [M. Duan and M. Urban, Phys. Rev. C 110, 065806 (2024)] independently without altering the properties of spin saturated matter. The parameters of the spin-dependent terms are determined by fitting the Landau parameters $G_{0}$ and $G_{0}^{\prime}$ in neutron matter and symmetric nuclear matter and the effective-mass splitting of up and down particles in spin polarized matter to the results of microscopic calculations. Using the new parametrizations, called Sky3s and Sky4s, the spin-related properties of nuclear matter are in good agreement with the microscopic results. As an application, we compute response functions and neutrino scattering rates of neutron-star matter with the new functionals having realistic effective masses and Landau parameters.
I Introduction
Skyrme effective interactions are easy to use due to their zero-range character. The energy density functional derived from Skyrme interactions has been widely used to study nuclear structure and neutron star properties (Vautherin and Brink, 1972; Beiner et al., 1975; Köhler, 1976; Chabanat et al., 1997; Goriely et al., 2010). To compute neutron star properties, interaction SLy4, constructed in 1998 (Chabanat et al., 1998), is very popular. Since 2010, a series of BSk interactions have also been developed for astrophysical applications (Goriely et al., 2010, 2013, 2016).
An important application of the nuclear energy density functional is to describe the transport properties of neutrinos in (proto-)neutron star and supernova matter. Neutrinos play a crucial role in explaining supernova explosions, neutron star mergers, proto-neutron star evolution, and neutron star cooling (Hirata et al., 1987; Mezzacappa, 2005; Janka, 2012; Burrows, 2013; Burrows and Lattimer, 1986; Mayle et al., 1987; Keil, 1994; Pascal et al., 2022; Yakovlev and Pethick, 2004). These astrophysical processes occur along with neutrino absorption and scattering. As many studies have done, such as the early literature Iwamoto and Pethick (1982) and the recent references Pastore et al. (2015); Oertel et al. (2020); Duan and Urban (2023), the neutrino rates (absorption and scattering rates) or the neutrino mean free path can be related to the response functions, including responses computed using the Skyrme interactions.
A couple of articles have studied neutrino mean free path using Skyrme energy density functional for pure neutron matter (PNM) (Pastore et al., 2012, 2014, 2015) and asymmetric nuclear matter (Davesne et al., 2014, 2019). But we recently reported that many Skyrme effective interactions predict an unrealistic density dependence of the effective masses, resulting, e.g., in the unphysical feature that the neutron Fermi velocity exceeds the speed of light at relatively low densities (Duan and Urban, 2023). Therefore new parametrizations were needed to solve this problem. We constructed two new Skyrme interactions, Sky3 and Sky4, solving the problem by adjusting the parameters to Brueckner-Hartree-Fock (BHF) results for the effective masses (Duan and Urban, 2024).
However, there is still a problem in these two interactions, namely, they predict a ferromagnetic instability at densities that may be realized in neutron stars (Duan and Urban, 2024). The reason is that the spin dependent terms of the functionals are expressed in terms of the same parameters as the spin-independent ones, and the fitting focuses on observables that depend only on the spin-independent terms with the exception of the spin-orbit parameter $W_{0}$. Apart from the unphysical instability, the spin dependent terms affect also the neutrino scattering rates which involve contributions of spin-0 (spin-independent) and spin-1 (spin-dependent) channels. The response functions of neutron-star and supernova matter, thus the neutrino rates, can be computed only if the Landau parameters in the spin-1 channel have the correct values. Therefore, the spin-dependent terms of the new Skyrme functionals should be constrained reasonably. This is possible if one interprets the Skyrme interaction as a density functional, in which the spin-independent and the spin-dependent terms can be adjusted independently Bender et al. (2002).
In this work, we first point out the problems of the spin-related properties of nuclear matter with the unconstrained Skyrme functionals Sky3 and Sky4, but also with other Skyrme functionals, in Sec. II. Then, we present the strategies to adjust the spin-dependent terms of the functionals by using results of BHF calculations in Sec. III. The Landau parameters in pure neutron matter and symmetric nuclear matter (SNM), the equation of state of spin-polarized neutron matter, and the instabilities in neutron star matter computed with the resulting new functionals Sky3s and Sky4s are also shown in Sec. III. Finally, we present some results of the full RPA (random phase approximation) response functions of neutron star matter and the corresponding neutrino scattering rates computed with Sky3s in Sec. IV, and summarize in Sec. V.
II Spin-related properties of nuclear matter with Skyrme functionals
The standard form of the Skyrme energy density functional consists of the kinetic-energy term, the zero-range term, the density-dependent term, the effective-mass term, the finite-range term, the spin-orbit term, the tensor coupling term, and the Coulomb energy density (Chabanat et al., 1997). The actual expressions can vary from one parametrization to another. For example, density-dependent generalizations of the non-local ($t_{1}$ and $t_{2}$) terms of Skyrme interactions have been added into the BSk functionals (Chamel et al., 2009). In interactions developed for astrophysical applications, the tensor coupling term is usually omitted (Goriely et al., 2010, 2013, 2016). The KIDS interactions have generalized the density dependent ($t_{3}$) term to several $t_{3i}$ terms with corresponding exponents $\alpha_{i}$ (Gil et al., 2019). In our previous work (Duan and Urban, 2024), the functionals Sky3 and Sky4 have been constructed combining the modification of the density-dependent term with the density-dependent generalizations of the $t_{1}$ and $t_{2}$ terms.
According to Refs. Bender et al. (2003); Duan and Urban (2023), the spin-dependent parts of the Skyrme energy density functional can be written using standard notations as
$\varepsilon_{\text{spin}}=\;$ $C_{0}^{s}\bm{\mathrm{s}}^{2}+C_{1}^{s}(\bm{\mathrm{s}}_{n}-\bm{\mathrm{s}}_{p})^{2}+C_{0}^{sT}(\bm{\mathrm{s}}\cdot\bm{\mathrm{T}}-\mathbbm{J}^{2})$
$+C_{1}^{sT}[(\bm{\mathrm{s}}_{n}-\bm{\mathrm{s}}_{p})\cdot(\bm{\mathrm{T}}_{n}-\bm{\mathrm{T}}_{p})-(\mathbbm{J}_{n}-\mathbbm{J}_{p})^{2}]$
$+C_{0}^{\Delta s}\bm{\mathrm{s}}\cdot\Delta\bm{\mathrm{s}}+C_{1}^{\Delta s}(\bm{\mathrm{s}}_{n}-\bm{\mathrm{s}}_{p})\cdot\Delta(\bm{\mathrm{s}}_{n}-\bm{\mathrm{s}}_{p})$
$+C_{0}^{\nabla\otimes s}(\bm{\mathrm{\nabla}}\otimes\bm{\mathrm{s}})^{2}+C_{1}^{\nabla\otimes s}[\bm{\mathrm{\nabla}}\otimes(\bm{\mathrm{s}}_{n}-\bm{\mathrm{s}}_{p})]^{2}$
$+C_{0}^{\nabla J}(\rho\bm{\mathrm{\nabla}}\cdot\bm{\mathrm{J}}+\bm{\mathrm{s}}\cdot\bm{\mathrm{\nabla}}\times\bm{\mathrm{j}})$
$+C_{1}^{\nabla J}[(\rho_{n}-\rho_{p})\bm{\mathrm{\nabla}}\cdot(\bm{\mathrm{J}}_{n}-\bm{\mathrm{J}}_{p})$
[…]
where (with the notation $y_{i}=t_{i}x_{i}$)
$C_{0}^{s}=\;$ ${-\frac{1}{8}}(t_{0}-2y_{0})-\sum_{i=1}^{3}\frac{1}{48}(t_{3i}-2y_{3i})\rho^{\alpha_{i}},$ (2a)
$C_{1}^{s}=\;$ ${-\frac{1}{8}}t_{0}-\sum_{i=1}^{3}\frac{1}{48}t_{3i}\rho^{\alpha_{i}},$ (2b)
$C_{0}^{sT}=\;$ $\frac{\eta_{J}}{16}[-(t_{1}-2y_{1})+(t_{2}+2y_{2})$
$-(t_{4}-2y_{4})\rho^{\beta}+(t_{5}+2y_{5})\rho^{\gamma}],$ (2c)
$C_{1}^{sT}=\;$ $\frac{\eta_{J}}{16}(-t_{1}+t_{2}-t_{4}\rho^{\beta}+t_{5}\rho^{\gamma}),$ (2d)
$C_{0}^{\Delta s}=\;$ $\frac{1}{32}[(t_{1}-2y_{1})+(t_{4}-2y_{4})\rho^{\beta}],$ (2e)
$C_{1}^{\Delta s}=\;$ $\frac{1}{32}(t_{1}+t_{4}\rho^{\beta}),$ (2f)
$C_{0}^{\nabla\otimes s}=\;$ ${-\frac{1}{64}}[(t_{1}-2y_{1})+(t_{2}+2y_{2})$
$+(t_{4}-2y_{4})\rho^{\beta}+(t_{5}+2y_{5})\rho^{\gamma}],$ (2g)
$C_{1}^{\nabla\otimes s}=\;$ ${-\frac{1}{64}}(t_{1}+t_{2}+t_{4}\rho^{\beta}+t_{5}\rho^{\gamma}),$ (2h)
$C_{0}^{\nabla J}=\;$ ${-\frac{3}{4}}W_{0},$ (2i)
$C_{1}^{\nabla J}=\;$ ${-\frac{1}{4}}W_{0}.$ (2j)
The parameter $\eta_{J}$, in the notation of Bender et al. (2003), indicates whether the contribution of the $t_{1}$ and $t_{2}$ terms to the spin-orbit field should be included ($\eta_{J}=1$) or not ($\eta_{J}=0$). For the relation between the spin current $\mathbbm{J}$ and the vector spin current $\bm{\mathrm{J}}$ see e.g. Ref. Lesinski et al. (2007).
Figure 1: The Landau parameters $G_{0}$ in pure neutron matter (top panel) and $G_{0}$ and $G_{0}^{\prime}$ in symmetric nuclear matter (middle and bottom panels). The results are computed with Skyrme interactions Sky3, Sky4 (Duan and Urban, 2024), SLy4 (Chabanat et al., 1998), BSk20 (Goriely et al., 2010), SLy5st (Margueron and Sagawa, 2009) and BSk17st (Margueron et al., 2009). The square symbols represent the microscopic BHF results of Refs. Zuo et al. (2003a, b).
In terms of $C_{0}^{s}$, $C_{1}^{s}$, $C_{0}^{sT}$, and $C_{1}^{sT}$, the Landau parameter $G_{0}$ in pure neutron matter can be written as
$G_{0}(\text{PNM})=2N_{0}[C_{0}^{s}+C_{1}^{s}+k_{F}^{2}(C_{0}^{sT}+C_{1}^{sT})]\,,$ (3)
while the Landau parameters $G_{0}$ and $G_{0}^{\prime}$ in symmetric nuclear matter can be written as
$G_{0}(\text{SNM})=4N_{0}(C_{0}^{s}+C_{0}^{sT}k_{F}^{2}),$ (4)
and
$G_{0}^{\prime}(\text{SNM})=4N_{0}(C_{1}^{s}+C_{1}^{sT}k_{F}^{2}),$ (5)
where $N_{0}=\frac{m^{*}k_{F}}{\pi^{2}\hbar^{2}}$, and $k_{F}=(3\pi^{2}\rho)^{1/3}$ for pure neutron matter and $k_{F}=(\frac{3}{2}\pi^{2}\rho)^{1/3}$ for symmetric nuclear matter.
The Landau parameters computed with Sky3 and Sky4 were already shown in Ref. Duan and Urban (2024). We display them again in Fig. 1 together with the results computed with other Skyrme interactions, such as SLy4 (Chabanat et al., 1998), BSk20 (Goriely et al., 2010), SLy5st (Margueron and Sagawa, 2009), and BSk17st (Margueron et al., 2009). SLy4 and BSk20 are, respectively, examples for standard and extended Skyrme interactions. SLy5st and BSk17st represent Skyrme interactions whose spin-dependent parts have been modified in an attempt to constrain them using BHF results (Margueron and Sagawa, 2009; Margueron et al., 2009). For comparison, the microscopic BHF results (Zuo et al., 2003a, b)  [^1] From Fig. 2 of Ref. Zuo et al. (2003) one can infer that there is a typo in Table 1 of Ref. Zuo et al. (2003a) and $G_{0}$ in pure neutron matter for $\rho=0.8\;\text{fm}^{-3}$ should read $3.52$ instead of $2.52$, but probably this value is still incorrect, cf. discussion below Eq. (13). are shown as square symbols. As we can see, the results computed with Sky3 and Sky4 are nearest to the BHF values at saturation density, but their evolutions with density are very different from the BHF results. Other Skyrme interactions present not only different evolutions but also different values at saturation density, compared with BHF results. We conclude that the fitting of the parameters of the considered interactions is not sufficient to make them predictive in the spin-1 channel. And even with the additional parameters introduced in SLy5st (Margueron and Sagawa, 2009) and BSk17st (Margueron et al., 2009) there is not enough freedom to reproduce the BHF results.
Figure 2: Equations of state of fully spin-polarized and unpolarized neutron matter computed with Sky3 and Sky4. The vertical red and steel-blue dotted lines represent the critical densities in pure neutron matter described using Sky3 and Sky4, respectively.
In the case of Sky3 and Sky4, the situation is worse because above some critical densities they give $G_{0}<-1$, violating the Landau stability criterion. To further illustrate this instability problem, we show in Fig. 2 the equations of state (EoS) of fully spin-polarized and unpolarized neutron matter computed with Sky3 and Sky4. For both Sky3 and Sky4, the fully polarized neutron matter becomes energetically favorable over the unpolarized one at densities above the respective critical densities indicated by the red and steel-blue dotted lines. The fundamental cause of this behaviour is the wrong values of the Landau parameters mentioned above. Otherwise, one would get a stiffer EoS for fully polarized than for unpolarized neutron matter. Therefore, the solution is to revise Sky3 and Sky4 to obtain realistic values of the Landau parameters.
III Parameters and nuclear-matter properties of the new Skyrme functionals
III.1 Determination of the spin-dependent terms of the new Skyrme functionals
From now on, we treat the functional as a general density functional instead of a functional derived from a density-dependent two-body force. According to Ref. Bender et al. (2002), the spin-independent and spin-dependent terms in the energy functional can be adjusted independently. Actually, the common choice $\eta_{J}=0$ (which we also adopted in Duan and Urban (2024) for Sky3 and Sky4) is already an example for this philosophy since the functional derived from the two-body force would correspond to $\eta_{J}=1$.
To distinguish them from the previous functionals Sky3 and Sky4, we will denote the new functionals with independently adjusted spin-dependent terms by Sky3s and Sky4s, respectively. First of all, analogously to the common choice $\eta_{J}=0$, we set
$C_{0}^{\nabla\otimes s}=C_{1}^{\nabla\otimes s}=C_{0}^{\Delta s}=C_{1}^{\Delta s}=0,$ (6)
since these coefficients are completely unconstrained by the observables (energies and radii of doubly-magic nuclei) to which we have fitted the parameters.
Figure 3: Fits to microscopic results of Bombaci et al. (2006); Bigdeli et al. (2009); Bigdeli and Bordbar (2024); Vidaña (2024) for $C^{sT}$ as functions of density in pure neutron matter (red) and symmetric nuclear matter (blue). The scattering of the microscopic results for fixed densities comes from deducing $C^{sT}$ at different values of the polarization.
We now focus on constraining the remaining coefficients $C_{0}^{s}$, $C_{1}^{s}$, $C_{0}^{sT}$, and $C_{1}^{sT}$. Notice that the choice $\eta_{J}=0$ implies that, in Sky3 and Sky4, we have $C_{0}^{sT}=C_{1}^{sT}=0$. This means that in spin-polarized matter, the spin up and down nucleons have the same effective mass, since
$\frac{m}{m^{*}_{\uparrow}}-\frac{m}{m^{*}_{\downarrow}}=\frac{4m}{\hbar^{2}}C^{sT}(\rho_{\uparrow}-\rho_{\downarrow})\,,$ (7)
where $C^{sT}=C_{0}^{sT}$ in the case of symmetric nuclear matter and $C^{sT}=C_{0}^{sT}+C_{1}^{sT}$ in the case of pure neutron matter. However, the BHF calculations of Ref. Bombaci et al. (2006) for spin-polarized neutron matter and calculations within the lowest-order constrained variational (LOCV) method for spin-polarized symmetric matter Bigdeli et al. (2009) found a splitting of the effective masses of spin up and down nucleons.
Since the spin-independent and spin-dependent terms in the energy functional can be adjusted independently (Bender et al., 2002), we will determine the $C^{sT}$ terms from these splittings rather than relating them to the parameters $t_{1}\dots y_{5}$ that were adjusted only to spin-independent observables. Assuming the same form of the density dependence as in Eq. (2) with $\beta=\gamma=\frac{1}{3}$ as in Sky3 and Sky4, we write $C^{sT}_{I}$ ($I=0,1$) as
$C^{sT}_{I}=C^{sT}_{I0}+C^{sT}_{I1}\rho^{\frac{1}{3}}.$ (8)
Then $C^{sT}_{00}$ and $C^{sT}_{01}$ can be obtained by fitting the splitting between the up and down nucleon effective masses for spin-polarized symmetric nuclear matter, while $C^{sT}_{00}+C^{sT}_{10}$ and $C^{sT}_{01}+C^{sT}_{11}$ can be obtained by fitting the splitting between the up and down neutron effective masses for spin-polarized neutron matter. To get information about the density dependence, the results published in Refs. Bombaci et al. (2006); Bigdeli et al. (2009) are not enough and we obtained more data at different densities directly from the authors of these papers Bigdeli and Bordbar (2024); Vidaña (2024). We show $C^{sT}$ as functions of density in pure neutron matter and symmetric nuclear matter in Fig. 3. As we can see, $C^{sT}_{I}$ cannot be considered constant.
Besides, as mentioned above, the $C_{I}^{s}$ and $C_{I}^{sT}$ coefficients can be related to the Landau parameters $G_{0}$ and $G_{0}^{\prime}$ in pure neutron matter and symmetric nuclear matter. Therefore, we are now concentrating on obtaining realistic values of the Landau parameters for our two functionals. As we did above for the $C^{sT}$ coefficients and in Ref. Duan and Urban (2024) for the effective masses, we will again use the results of microscopic BHF calculations for this purpose.
We now rewrite $C_{0}^{s}$ and $C_{1}^{s}$ as
$C_{0}^{s}=-\frac{1}{8}(t_{0}-2y_{0})-\sum_{i=1}^{3}\frac{1}{48}(t_{3i}^{s}-2y_{3i}^{s})\rho^{\alpha_{i}},$ (9)
and
$C_{1}^{s}=-\frac{1}{8}t_{0}-\sum_{i=1}^{3}\frac{1}{48}t_{3i}^{s}\rho^{\alpha_{i}}.$ (10)
This means $t_{3i}^{s}$ and $y_{3i}^{s}$ are different from $t_{3i}$ and $y_{3i}$ in the spin-independent terms, but $t_{0}$ and $y_{0}$ are kept because the interaction is dominated by two-body interaction at very low densities. The $C^{s}_{I}$ coefficients ($I=0,1)$ can also be expressed as
$C^{s}_{I}=C^{s}_{I0}+C^{s}_{I1}\rho^{\frac{1}{3}}+C^{s}_{I2}\rho^{\frac{2}{3}}+C^{s}_{I3}\rho,$ (11)
with
$C^{s}_{00}$ $=-\frac{1}{8}(t_{0}-2y_{0}),$ (12a)
$C^{s}_{0i}$ $=-\frac{1}{48}(t_{3i}^{s}-2y_{3i}^{s})$ $(i=1,2,3),$ (12b)
$C^{s}_{10}$ $=-\frac{1}{8}t_{0},$ (12c)
$C^{s}_{1i}$ $=-\frac{1}{48}t_{3i}^{s}$ $(i=1,2,3).$ (12d)
Hence, only $C^{s}_{00}$ and $C^{s}_{10}$ are defined by the values of $t_{0}$ and $y_{0}$ of Sky3 or Sky4, respectively. Because we suspect that there was a mistake in the calculation of $G_{0}$ in pure neutron matter in Refs.Zuo et al. (2003a, 2003) (see below), we determine $C^{s}_{I1}\dots C^{s}_{I3}$ by fitting only the Landau parameters $G_{0}$ and $G_{0}^{\prime}$ in symmetric nuclear matter to those computed using the microscopic BHF theory in Ref. Zuo et al. (2003b). According to Zuo et al. (2003), $G_{0}$ in pure neutron matter is computed from the spin susceptibility $\chi$ as
$G_{0}=\frac{m^{*}}{m}\left(\frac{\chi}{\chi_{F}}\right)^{-1}-1,$ (13)
where $\chi_{F}$ is the spin susceptibility of an ideal gas. By analyzing the different results for $G_{0}$ and $\chi$ shown in Ref. Zuo et al. (2003), one finds that in the calculation with three-body force, the factor $\frac{m^{*}}{m}$ in Eq. (13) was by mistake taken to be the same as in the calculation without three-body force. We can compute corrected values of $G_{0}$ from Eq. (13) using the results for $\tfrac{\chi}{\chi_{F}}$ given in Zuo et al. (2003) and for $\tfrac{m^{*}}{m}$ from Baldo et al. (2014) (which are similar to the Sky3 and Sky4 ones). We use only the last point (at $\rho=0.8\;\text{fm}^{-3}$) of the corrected $G_{0}$ to roughly guide the high-density behavior of the fits. The results of these fits are listed in Table 1.
Table 1: Values of the new parameters $C^{sT}_{Ii}$, $C^{s}_{Ii}$, and $W_{0}$ for Sky3s and Sky4s. All other parameters are the same as in Sky3 and Sky4, respectively.
Sky3s Sky4s
$C^{sT}_{00}$ (MeV fm5) -43.6 -43.6
$C^{sT}_{01}$ (MeV fm6) 42.2 42.2
$C^{sT}_{10}$ (MeV fm5) -15.0 -15.0
$C^{sT}_{11}$ (MeV fm6) 33.0 33.0
$C^{s}_{00}$ (MeV fm3) 338.7 314.3
$C^{s}_{01}$ (MeV fm4) -1059.4 -945.7
$C^{s}_{02}$ (MeV fm5) 1414.6 1239.9
$C^{s}_{03}$ (MeV fm6) -406.7 -318.0
$C^{s}_{10}$ (MeV fm3) 225.1 242.9
$C^{s}_{11}$ (MeV fm4) -385.3 -468.3
$C^{s}_{12}$ (MeV fm5) 336.7 464.2
$C^{s}_{13}$ (MeV fm6) -32.7 -97.4
$W_{0}$ (MeV fm5) 117.4 114.1
Since we have now non-vanishing $C_{I}^{sT}$, the spin-orbit splitting in finite nuclei is modified, because the spin-orbit field $U_{ls,q}$ (in the notation of Reinhard (1991), $q=n,p$) acquires an additional term
$U_{ls,q}^{\text{(add)}}=(C^{sT}_{1}-C^{sT}_{0})J-2C^{sT}_{1}J_{q}\,.$ (14)
Because of the density dependence of $C^{sT}_{I}$, there are also additional terms in the mean-field potential,
$U^{\text{(add)}}=\frac{1}{2}(C_{1}^{sT\,\prime}-C_{0}^{sT\,\prime})J^{2}-C_{1}^{sT\,\prime}(J_{n}^{2}+J_{p}^{2})\,,$ (15)
(with $C_{I}^{sT\,\prime}=dC_{I}^{sT}/d\rho=C_{I1}^{sT}\rho^{-2/3}/3$), and in the rearrangement energy,
$E_{\text{rearr}}^{\text{(add)}}=-\int\!d^{3}r\,\frac{1}{2}\,\rho\,U^{\text{add}}.$ (16)
In order to still get good binding energies and radii, we therefore have to refit the spin-orbit parameter $W_{0}$. Refitting only $W_{0}$ and keeping the other parameters as in Ref. Duan and Urban (2024) is sufficient because after the refitting, the $\chi^{2}$ of binding energies and radii as defined in Duan and Urban (2024) is 4.48 for Sky3s and 5.55 for Sky4s, which is almost as good as the $\chi^{2}$ of 4.08 and 5.07 obtained for Sky3 and Sky4. The new values of $W_{0}$ can also be found in Table 1. We tried refitting also the other parameters, but the resulting improvement of $\chi^{2}$ was negligible.
III.2 The lowest-order Landau parameters in nuclear matter and spin-polarized neutron matter EoS
Figure 4: The Landau parameters $G_{0}$ in pure neutron matter (upper panel) and $G_{0}$ and $G_{0}^{\prime}$ in symmetric nuclear matter (lower panel) computed with Sky3s and Sky4s. The BHF results from Refs. Zuo et al. (2003a, b) are also shown.
Figure 5: EoS in fully spin-polarized and unpolarized neutron matter for Sky3s and Sky4s. The results from Refs. Vidaña et al. (2016); Tews and Schwenk (2020) for polarized neutron matter are also shown for comparison.
In this section, we present some nuclear matter properties described by the two new interactions with the new parameters as explained above.
In Fig. 4, we show the Landau parameters $G_{0}$ in pure neutron matter (upper panel) and $G_{0}$ and $G_{0}^{\prime}$ in symmetric nuclear matter (lower panel). We also present the results computed using the microscopic BHF theory in Refs. Zuo et al. (2003a, b) for comparison. All values are larger than 0 ($>-1$). This means that there is no ferromagnetic instability in spin-1 channels. The $G_{0}$ and $G_{0}^{\prime}$ in symmetric nuclear matter at densities $0.17-0.45\;\text{fm}^{-3}$ agree within 10% with the BHF results (only for the points at the lowest density, $0.08\;\text{fm}^{-3}$, the deviations are slightly larger). At low densities, the agreement is not as good as in symmetric matter but still satisfactory. Besides, $G_{0}^{\prime}\approx 1.20$ for Sky3s and $G_{0}^{\prime}\approx 1.19$ Sky4s at saturation density. These values also agree well with the result $G_{0}^{\prime}\approx 1.2$ of Ref. Bender et al. (2002) obtained from Gamow-Teller resonances in finite nuclei.
In Fig. 5, we show the EoS in fully spin-polarized and unpolarized neutron matter for Sky3s and Sky4s. The microscopic BHF results (Vidaña et al., 2016) and the AFDMC calculation with the N2LO $V_{E1}$ of Ref. Tews and Schwenk (2020) are also shown for comparison. From the comparison between the EoS of polarized and unpolarized neutron matter, we can see that spin-polarized neutron matter always has higher energy than unpolarized one at all densities for both Sky3s and Sky4s. This is only possible because now also the energy of polarized matter increases strongly at high density. This agrees qualitatively with the results in Ref. Vidaña et al. (2016), although our energies are larger than theirs (the grey curve in Fig. 5).
We can see that Sky3s and Sky4s give a slightly stiffer EoS in spin-polarized neutron matter than BHF and AFDMC calculations, even at densities below $2\rho_{0}$. This does not affect the use of our new interactions in dealing with neutron star matter. On the one hand, the strong spin polarization does not exist in nature. On the other hand, getting realistic values for Landau parameters is more important and relevant, in particular in the context of neutrino rates.
III.3 Instabilities of neutron star matter described by the two new interactions
There can be singularities at zero energy transfer in response functions of nuclear matter. They indicate the appearance of instabilities. The positions of instabilities should be determined to see whether the new interactions can describe neutron star matter. We will compute and show the results for $\beta$-stable neutron star matter in this section.
In the channel $S=0$ ($S$ is the spin), the so-called spinodal instability (Ducoin et al., 2008) is physical. This spinodal instability represents the liquid-gas phase transition, which has been observed in symmetric nuclear matter in experimental studies (Suraud et al., 1989). In the case of neutron star matter, it corresponds to the core-crust transition. This transition can strongly affect the transport process of neutrinos in hot and dense nuclear matter (Margueron and Sagawa, 2009; Reddy et al., 1999; Navarro et al., 1999). In other words, this transition can affect the supernova explosion and proto-neutron star evolution because neutrinos play a crucial role in these astrophysical processes.
However, other instabilities are unphysical. For small momentum transfer $q\approx 0$, we can also use the Landau stability criterion as in the cases of pure neutron matter and symmetric nuclear matter. In general, except for the spinodal instability, the instabilities of neutron star matter (zero temperature) occur at densities that are near the critical densities of pure neutron matter and symmetric nuclear matter (Davesne et al., 2014; Pastore et al., 2015).
For different values of the baryon number density $\rho$ of neutron-star matter and of the momentum transfer $q$, we can find the onset of the instability from the condition (Pastore et al., 2014)
$\frac{1}{\Pi^{(S,M,I)}_{\text{RPA}}(q,\omega=0)}=0,$ (17)
($M$ is the projection of spin along the direction of $\bm{\mathrm{q}}$, $I$ is the isospin) to detect poles in the RPA response function $\Pi_{\text{RPA}}$ after determining the proton fraction $Y_{p}$ through the $\beta$-equilibrium condition at zero temperature.
Figure 6: The onset of spin-1 instabilities in the $q-\rho$ plane for Sky3s and Sky4s for neutron star matter in $\beta$ equilibrium, compared to those of the previous unconstrained parametrizations Sky3 and Sky4.
We show the positions of the onset of the instabilities for Sky3s and Sky4s in the $q-\rho$ plane in Fig. 6. The results for Sky3 and Sky4 are also shown for comparison. We show only the results of spin-1 channels since the results of spin-0 channels have been shown in Ref. Duan and Urban (2024).
In the case of the unconstrained interactions Sky3 and Sky4, we can see from Fig. 6 that the instabilities can be found at densities that are lower than the maximum central density of a neutron star ($\rho_{\text{max}}=0.9248\;\text{fm}^{-3}$ for Sky3, $\rho_{\text{max}}=1.1264\;\text{fm}^{-3}$ for Sky4 (Duan and Urban, 2024)) even at lower momentum transfers. Also, at higher momentum transfers, Sky3 and Sky4 predict instabilities of neutron star matter even when the baryon number density is lower than saturation density $\rho_{0}$. This is not acceptable.
In the case of the new constrained interactions Sky3s and Sky4s, we can see from Fig. 6 that the unphysical ferromagnetic instabilities in spin-1 channels can be found in the full RPA responses of neutron star matter only at high densities and high momentum transfers. At densities near the maximum central densities of neutron stars predicted by Sky3s and Sky4s, there is no unphysical ferromagnetic instability until the momentum transfer is higher than about $2\;\text{fm}^{-1}$. At such high momentum transfers, Skyrme interactions should not be used. But these momentum transfers do not appear in neutrino scattering.
IV Response functions and neutrino scattering rates in neutron-star matter
Figure 7: Spin-1 channel response functions computed with Sky3s. Neutron star matter is at zero temperature and $\beta$ equilibrium. $\rho=0.25\;\text{fm}^{-3}$, $Y_{p}=0.0646$, $q=0.5\;\text{fm}^{-1}$. The results computed with Sky3 are also shown for comparison. Upper panel: isoscalar response functions ($I=0$); lower panel: isovector response functions ($I=1$).
Figure 8: Differential neutrino scattering rates as a function of the final neutrino energy $E_{\nu}^{\prime}$ with the fixed initial neutrino energy $E_{\nu}=3T$ for Sky3s (upper panel) and the corresponding evolution of the average cosine of the scattering angle (lower panel). $T=10\;\text{MeV}$, $\rho=0.25\;\text{fm}^{-3}$, $Y_{p}=0.0646$. The red dotted lines correspond to the initial neutrino energy. The results computed with Sky3 are also shown for comparison.
In this section, we show examples for full RPA response functions for $S=1$ channels and neutrino scattering rates computed using Sky3s.
Pastore et al. presented a detailed review of response functions computed using Skyrme energy density functional (Pastore et al., 2015). Another equivalent method to compute response functions was proposed in Ref. Urban and Ramanan (2020), which has been generalized for use in asymmetric nuclear matter and finite energy transfer (Duan and Urban, 2023). Here, we adopt the method described in Ref. Duan and Urban (2023) to compute the RPA response functions of neutron star matter. Following the steps explained in Section II B of Ref. Duan and Urban (2023), the expressions of RPA response functions in asymmetric nuclear matter can be obtained.
In general, the dynamical structure factors can be related to the RPA response functions as
$S^{(S,M,I)}(\bm{\mathrm{q}},\omega)=-\frac{1}{\pi}\frac{1}{1-e^{-\omega/T}}\operatorname{Im}\Pi^{(S,M,I)}_{\text{RPA}}(\bm{\mathrm{q}},\omega).$ (18)
We show the spin-1 dynamical structure factors $S(q,\omega)$ for Sky3s in neutron star matter at zero temperature under the condition of $\beta$ equilibrium in Fig. 7. The upper and lower panels show the isoscalar ($I=0$) and isovector ($I=1$) responses, respectively, for baryon number density $\rho=0.25\;\text{fm}^{-3}$, proton fraction $Y_{p}=0.0646$, and momentum transfer $q=0.5\;\text{fm}^{-1}$. The red arrows represent the positions of the zero-sound modes for Sky3s. The results computed with Sky3 are also shown for comparison. As we can see, the responses computed with Sky3 are larger than those computed with Sky3s. Besides, the zero-sound modes lie in the particle-hole continuum for Sky3. These differences are due to the fact that the unconstrained Landau parameters of Sky3 are not repulsive enough, which can affect computations of neutrino rates in (proto-)neutron star and supernova matter.
As mentioned in the introduction, neutrinos play a crucial role in explaining some astrophysical phenomena. The interactions between neutrinos and nuclear matter should not be ignored. In particular, neutrino scattering is as important as neutrino absorption.
We compute the differential neutrino scattering rate and average cosine of the scattering angle using the method described in Section IV C of Ref. Duan and Urban (2023) and show the results in Fig. 8. To include the contribution of the zero-sound modes, we compute at temperature $T=10\;\text{MeV}$ (so that the zero-sound mode gets a finite width) but keep the baryon number density and proton fraction as in Fig. 7. The results are plotted as functions of the final neutrino energy $E_{\nu}^{\prime}$, for the initial neutrino energy $E_{\nu}=3T$ (red dotted lines in Fig. 8).
As anticipated, Sky3s gives different scattering rate and average scattering angle from Sky3 at the same $E_{\nu}^{\prime}$. For most values of $E_{\nu}^{\prime}$, Sky3s gives smaller neutrino scattering rates and average scattering angles (larger average cosine of the scattering angle) than Sky3. These can affect simulations of proto-neutron star evolution and supernova explosion.
V Conclusion
This work aims to determine the spin-dependent terms of the new Skyrme functionals introduced in Ref. Duan and Urban (2024) to obtain realistic values for the Landau parameters $G_{0}$ and $G^{\prime}_{0}$, and thus a reasonable description of the spin-related properties of nuclear matter.
Since it is not easy to get precise information on spin dependent terms from experiment, especially for neutron-rich matter at high density, we have fitted the values of the Landau parameters and the effective-mass splittings in spin polarized matter for Sky3s and Sky4s to microscopic BHF results. In this way, most of the spin-dependent parts of the two new interactions have been constrained. This is in the same spirit as what was done in Ref. Duan and Urban (2024) for the spin-independent terms by fitting BHF results for the neutron-matter EoS and the effective masses.
However, while the spin-independent gradient terms were constrained in Ref. Duan and Urban (2024) from the fit to finite nuclei, we have no information about the spin-dependent gradient terms and we chose to set them equal to zero. These terms of the functional should be refined in future studies.
Now, the response functions can be computed with the new Skyrme functionals having realistic effective masses and Landau parameters. Since the typical momentum transfers in neutrino scattering processes in (proto-)neutron stars are not very high, this is sufficient to get reasonable neutrino rates, while the still undetermined spin-dependent gradient terms contribute only at higher momentum transfers.
Concerning uniform matter properties, the new functionals do of course not provide more predictivity than the BHF calculations to which they were fitted. But they are much easier to use, in particular, for the computation of the EoS and of the composition of neutron-star matter in $\beta$ equilibrium, for finite temperature calculations, response functions, and neutrino rates. Furthermore, since they were also fitted to finite nuclei, they should also be suitable for the description of the neutron-star crust. Therefore, the two new constrained interactions Sky3s and Sky4s will be useful to continue towards the objective of Ref. Duan and Urban (2023) to compute neutrino rates for astrophysical simulations. More generally, they can be used in all kind of future studies in nuclear astrophysics where simple Skyrme energy functionals are needed because other approaches are computationally too costly.
It should be mentioned, however, that the old BHF results for the Landau parameters of Refs. Zuo et al. (2003a, b, 2003) that we used for our fits have some problems. For instance, the results for $G_{0}$ in symmetric nuclear matter shown in Refs. Zuo et al. (2003b) do not agree with those of Ref. Zuo et al. (2003). And in addition to the problem discussed after Eq. (13), it should be mentioned that using those values for $G_{0}$ in pure neutron matter, the energy of fully spin polarized neutron matter is too high compared to the results of Tews and Schwenk (2020); Vidaña et al. (2016) as seen in Fig. 5. Therefore, it would be important that the question of spin Landau parameters be revisited by some new microscopic calculations.
Acknowledgements.
We thank M. Bigdeli, G. H. Bordbar, and I. Vidaña for sending us their unpublished results for the effective masses in spin-polarized matter. Mingya Duan is grateful for the support of the China Scholarship Council (CSC No. 202006660002).